# PlainQAFact: Automatic Factuality Evaluation Metric for Biomedical Plain Language Summaries Generation

### v1

@@ Abstract
Hallucinated outputs from language models pose risks in the medical domain, especially for lay audiences making health-related decisions. Existing factuality evaluation methods, such as entailment- and question-answering-based (QA), struggle with plain language summary (PLS) generation due to elaborative explanation phenomenon, which introduces external content (e.g., definitions, background, examples) absent from the source document to enhance comprehension. To address this, we introduce PlainQAFact, a framework trained on a fine-grained, human-annotated dataset PlainFact, to evaluate the factuality of both source-simplified and elaboratively explained sentences. PlainQAFact first classifies factuality type and then assesses factuality using a retrieval-augmented QA-based scoring method. Our approach is lightweight and computationally efficient. Empirical results show that existing factuality metrics fail to effectively evaluate factuality in PLS, especially for elaborative explanations, whereas PlainQAFact achieves state-of-the-art performance. We further analyze its effectiveness across external knowledge sources, answer extraction strategies, overlap measures, and document granularity levels, refining its overall factuality assessment [^1] PlainQAFact and PlainFact are available at: https://github.com/zhiwenyou103/PlainQAFact.
PlainQAFact: Automatic Factuality Evaluation Metric for
Biomedical Plain Language Summaries Generation
Zhiwen You  Yue Guo University of Illinois Urbana-Champaign {zhiweny2, yueg}@illinois.edu
Figure 1: Illustration of our PlainQAFact framework. The inputs are PLSs and scientific abstracts. Our evaluation framework consists of a pre-trained classifier for factuality type categorization, retrieval of external domain knowledge for elaborative explanation sentences, and a local LLM-based question-answering system.
@@ 1 Introduction
Communicating biomedical scientific knowledge in plain language is essential for improving health information accessibility Stoll et al. (2022); Kuehne and Olden (2015). While large language models (LLMs) have made significant progress in generating plain language summaries (PLSs) Ondov et al. (2022); Guo et al. (2023); Goldsack et al. (2023); You et al. (2024b), the issue of hallucination remains unresolved Cao et al. (2022); Guo et al. (2024b).
Factuality evaluation for PLS is particularly challenging due to the presence of elaborative explanations, where external content such as definitions, background information, and examples is added to enhance comprehension but is absent from the original scientific abstract (source) Guo et al. (2021); Srikanth and Li (2021); Joseph et al. (2024). This creates two key issues. First, commonly used factuality metrics can effectively verify content supported by the source but fail to assess factual consistency of added information Guo et al. (2024a). Second, existing datasets for evaluating PLS factuality are limited in both data and annotation quality. FactPICO Joseph et al. (2024), for instance, focuses on LLM-generated summaries and annotates only newly added information, overlooking potential hallucinations from source simplifications. This limitation prevents a comprehensive evaluation of factual consistency across the entire PLS.
To address theses challenges, we hypothesize that improving factuality evaluation in elaborative explanations enhances overall PLS factuality assessment. To test this, we introduce the first granular factuality dataset for PLS, with expert annotations for factuality type (elaborative explanation vs. source simplification), functional role (e.g., definition, background, example), and sentence-alignment between human-authored PLS and corresponding scientific abstracts. These fine-grained annotations establish a benchmark for advancing PLS generation and factuality evaluation methods.
Building on this, we propose a dual-stage QA-based factuality evaluation framework PlainQAFact, that independently scores each PLS sentence (see Figure 1). First, a pre-trained classifier predicts whether a PLS sentence is a source simplification or an elaborative explanation. For elaborative explanations, dense retrieval via MedCPT Jin et al. (2023) retrieves relevant external information, for verification beyond the source abstract. The abstract and retrieved content are then concatenated as references for QA-based factuality evaluation. For source simplifications, the abstract alone serves as the reference for evaluation. Finally, sentence-level evaluation scores are averaged to produce a final factuality score for the PLS. Experiments on PlainFact and the CELLS dataset Guo et al. (2024b) demonstrate that PlainQAFact significantly enhances PLS factuality assessment, especially for elaborative explanations. By enabling more accurate factuality assessment in PLS, our approach enhances robustness and supports a foundation for enhancing open-ended PLS generation.
Our main contributions are as follows:
  • PlainFact: the first expert-annotated benchmark for granular factuality evaluation in PLS, covering factuality type, functional role, and sentence alignment (§3).
  • PlainQAFact: a factuality evaluation framework integrates sentence type classification with retrieval-augmented QA, leveraging expert-domain resources (§4).
  • PlainQAFact outperforms existing factuality metrics in PLS evaluation, particularly in handling elaborative explanations (§5, 6).
@@ 2 Related Work
The primary approach for evaluating PLS generation combines automated metrics with human evaluation (Jain et al., 2021; Ondov et al., 2022). While human assessment provides a thorough analysis Hardy et al. (2019), its high cost and time demands make it impractical for large-scale datasets. Evaluating factuality in biomedical PLS is particularly challenging, as it requires domain expertise. Entailment- Lee et al. (2022); Laban et al. (2022), similarity- Wan and Bansal (2022); Ye et al. (2024), model- Zha et al. (2023) and QA-based Deutsch et al. (2021); Fabbri et al. (2022b) metrics are commonly used for factuality assessment but rely heavily on high-quality reference summaries, which are often unavailable or difficult to obtain for PLS. Recent advancements in prompt-based evaluation show promise (Luo et al., 2023); however, their sensitivity to factuality perturbations in PLS remains limited Guo et al. (2024a). To address these limitations, we propose a reference-free solution for factuality evaluation of PLS that effectively assesses factual consistency with external information retrieval to augment the reference summary.
@@
Retrieval-augmented methods enhance text generation by extracting relevant information from external sources to supplement input queries Lewis et al. (2020). These methods have been shown to be effective in open-domain QA Ren et al. (2023); Mao et al. (2020), knowledge-based QA Kim (2025), and multi-step reasoning Gao et al. (2023); Tang and Yang (2024). In the context of PLS, retrieval from structured knowledge bases (KBs) has been shown to improve factual accuracy compared to language models alone Guo et al. (2024b). However, retrieval-augmented approaches have not been extensively explored for factuality evaluation in PLS, despite their potential for addressing elaborative explanations. In this work, we investigate retrieval-augmented QA to enhance PLS factuality assessment while also examining its limitations.
3 PlainFact Benchmark
To develop a high-quality PLS factuality benchmark, we select a subset of the largest human-authored PLS dataset (§3.1) and hire domain experts to provide fine-grained annotations (§3.2).
3.1 Human-Authored PLS Dataset
Rather than relying on LLM-generated PLSs, we construct our benchmark using human-authored summaries. CELLS Guo et al. (2024b), the largest parallel corpus of scientific abstracts and their corresponding PLSs, is written by the original authors and sourced from 12 biomedical journals. We primarily select data from the Cochrane Database of Systematic Reviews (CDSR) [^2] http:///www.cochranelibrary.com within CELLS, as CDSR contains systematic reviews that support evidence-based medical decision-making across healthcare domains Murad et al. (2016). Since systematic reviews represent the highest level of scientific evidence, this selection enhances the factual rigor of our dataset. To ensure readability, we filter the 200 most readable PLSs based on average scores from three standard readability metrics: Flesch-Kincaid Grade Level (FKGL) Kincaid (1975), Dale-Chall Readability Score (DCRS) Dale and Chall (1948), and Coleman-Liau Index (CLI) Coleman (1975).
Explanation Simplification
@@ 3.1 Human-Authored PLS Dataset
Chall $\downarrow$ 11.3 11.6
Coleman-Liau $\downarrow$ 13.5 13.9
@@
Table 1: Overview of the PlainFact benchmark. Medical experts annotated 200 pairs of PLSs and their corresponding scientific abstracts from the CELLS dataset, categorizing each PLS sentence as either a source simplification or an elaborative explanation related to the abstract.
3.2 Expert Annotation
@@
Since each PLS-abstract pair is authored by the same individual, we assume all information to be factual. The annotation aims to capture how PLS sentences relate to their scientific abstracts. Annotators analyzed each PLS sentence across three dimensions: (1) Factuality type: identify whether a sentence is a source simplification (derived from the abstract) or an elaborative explanation (introducing new content); (2) Functional role: categorize the sentence as background, definition, example, method/result, or other; and (3) Sentence alignment: map each PLS sentence to its corresponding sentence(s) in the scientific abstract. Details of annotation guidelines are in App. A.
@@
Annotations were conducted by four independent annotators, each with at least a bachelor’s degree in biomedical sciences and prior fact-checking experience. Annotators were recruited via Upwork and compensated from \$15 to \$20 per hour. Each PLS-abstract pair was annotated by two independent annotators, with disagreements resolved by a third. Inter-rater agreement, measured by Cohen’s Kappa for factuality type, functional role, and sentence alignment are 0.43, 0.60, and 0.55 respectively, indicated moderate agreement for all tasks Artstein and Poesio (2008).
@@
Table 1 summarizes the dataset characteristics. Notably, 44% of PLS sentences contain elaborative explanations, highlighting their role in enhancing PLS readability. This underscores the need for factuality evaluation methods that account for such phenomena. Moreover, elaborative explanations include more background, definitions, and examples than source simplifications. For annotation examples, see App. B.
4 PlainQAFact Framework
@@
PlainQAFact conducts fine-grained factuality evaluation for PLS by assessing each sentence individually and retrieving tailored external knowledge within a retrieval-augmented QA framework. Figure 1 provides an overview of its three key components: sentence-level classification (§4.1), domain knowledge retrieval (§4.2), and dual-stage QA-based factuality evaluation (§4.3).
4.1 Learned Factuality Type Classifier
@@
As elaborative explanations are prevalent in PLS generation and existing metrics struggle to capture added information Guo et al. (2024a), we fine-tune a classifier to categorize factuality types. Using factuality type annotations (source simplification vs. elaborative explanation) in PlainFact, we fine-tune a classifier with PubMedBERT-base model [^3] https://huggingface.co/microsoft/BiomedNLP-BiomedBERT-base-uncased-abstract-fulltext. The dataset is split 8:1:1 for training, validation, and testing. The classifier is fine-tuned with early stopping, stopping at the third epoch based on validation performance. Additionally, we compare the learned classifier with a zero-shot GPT-4o OpenAI (2024) classifier (details in App. C) to explore extending our factuality evaluation framework to domains and tasks lacking human-annotated data.
4.2 Domain Knowledge Retrieval
@@
Retrieval-augmented methods have proven effective for explanation generation Guo et al. (2024b); Xu et al. (2024), making them a natural fit for evaluating elaborative explanations. Since our dataset is in the medical domain, we adopt MedCPT Jin et al. (2023), a retriever pretrained on large-scale PubMed search logs to generate biomedical text embeddings. For external resources, we incorporate StatPearls Jin et al. (2021) for clinical decision support and medical textbooks Xiong et al. (2024) for domain-specific knowledge.
4.3 QA Evaluation Components
@@
QA-based metrics have outperformed other factuality metrics Guo et al. (2024a) in PLS and align more closely with human annotations Fabbri et al. (2022a); Deutsch et al. (2021). Therefore, we adopt a QA-based approach in our framework to verify factuality, incorporating answer extraction, question generation, question filtering, question answering, and answer overlap calculation.
@@
The first step in QA-based factuality evaluation is to extract answer entities (keyphrases) from the PLS as gold answer, then verify factuality by comparing them with answers generated by a QA model for the same questions. If the generated answer is correct or relevant, the summary is considered factual. To compare answer extraction strategies, we use noun phrase chunks (NP Chunks) as the baseline, and evaluate LLM-based extractors, including Llama 3.1 8B Instruct (Llama 3.1)Meta (2024), GPT-4o OpenAI (2024), and PromptRank Kong et al. (2023). Prompts for Llama and GPT-4o are in App. C. PromptRank is a keyphrase generation method based on the T5 model. The number of extracted entities is set to 50 and the T5-large is used for better performance.
@@
Given an input document (PLS), the QG model generates questions based on the extracted answers and their context. Following previous studies Deutsch et al. (2021); Fabbri et al. (2022b) [^4] These studies use allennlp, an open-source NLP research library built on PyTorch, which is no longer actively maintained., we fine-tune BART-large model on standard QG datasets, including SQuAD Rajpurkar et al. (2016) and QA2D Demszky et al. (2018), for use as the QG model in our framework.
@@
To prevent unanswerable questions from impacting the final factuality score, we follow QAFactEval Fabbri et al. (2022a) and remove the unanswerable questions using a pre-trained Electra-large model Clark et al. . During QF, the QA model receives only the PLS and its corresponding questions. In the subsequent QA stage, answers are extracted from the source for answer overlap evaluation, while QF only determines whether the questions can be answered by the PLS.
@@
The QA model extracts answers to filtered questions from the source document. To prevent hallucinated output, we use an extractive QA model, a pre-trained Electra-large, which was the best performing QA model in QAFactEval Fabbri et al. (2022a).
@@
We evaluate the alignment between the gold and generated answers from the QA model using the LERC score Chen et al. (2020) and BERTScore Zhang et al. . The final step in PlainQAFact is weighting the scores for source simplification and elaborative explanation sentences. Specifically, the simplification score $s$ is computed using only abstracts as the source in QA, while the explanation score $e$ incorporates both abstracts and retrieved knowledge.
@@
where $n_{s}$ is the number of simplification sentences, and $n_{e}$ is the number labeled as explanation.

### v2

# PlainQAFact: Retrieval-augmented Factual Consistency Evaluation Metric for Biomedical Plain Language Summarization

@@ Abstract
Hallucinated outputs from large language models (LLMs) pose risks in the medical domain, especially for lay audiences making health-related decisions. Existing automatic factual consistency evaluation methods, such as entailment- and question-answering (QA) -based, struggle with plain language summarization (PLS) due to elaborative explanation phenomenon, which introduces external content (e.g., definitions, background, examples) absent from the scientific abstract to enhance comprehension. To address this, we introduce PlainQAFact, an automatic factual consistency evaluation metric trained on a fine-grained, human-annotated dataset PlainFact, for evaluating factual consistency of both source-simplified and elaborately explained sentences. PlainQAFact first classifies sentence type, then applies a retrieval-augmented QA scoring method. Empirical results show that existing evaluation metrics fail to evaluate the factual consistency in PLS, especially for elaborative explanations, whereas PlainQAFact consistently outperforms them across all evaluation settings. We further analyze PlainQAFact’s effectiveness across external knowledge sources, answer extraction strategies, answer overlap measures, and document granularity levels, refining its overall factual consistency assessment. Taken together, our work presents the first evaluation metric designed for PLS factual consistency evaluation, providing the community with both a robust benchmark and a practical tool to advance reliable and safe plain language communication in the medical domain. PlainQAFact and PlainFact are available at: https://github.com/zhiwenyou103/PlainQAFact.
keywords:
Plain language summarization , Factual consistency evaluation , Retrieval-augmented generation , Hallucination , Large language models
[…]
[label1]organization=School of Information Sciences,
University of Illinois Urbana-Champaign,city=
Champaign, postcode=61820, state=IL, country=United States
@@ 1 Introduction
Communicating biomedical scientific knowledge in plain language is essential for improving health information accessibility and health literacy [1, 2]. Recent advances in large language models (LLMs) have made significant progress in plain language summarization (PLS) of biomedical texts [3, 4, 5, 6]. However, ensuring the factual consistency of these summaries remains a major challenge. A key source of inconsistency stems from elaborative explanations: content such as definitions, background information, and examples that enhance comprehension but are not explicitly present in the original scientific abstracts (i.e., source) [7, 8, 9]. While such elaborations are critical for effective communication, they introduce external content that cannot be directly verified against the source, complicating automatic factual consistency evaluation.
Factual consistency in PLS is typically assessed through a combination of human evaluation and automated metrics [3, 10]. While human evaluation is reliable [11], it is costly and difficult to scale, particularly in biomedical contexts where domain expertise is required. Commonly used factuality metrics can effectively verify content supported by the source but fail to assess factual consistency of added information [12]. However, these metrics depend heavily on high-quality reference summaries, which are often unavailable in plain language summaries. Recent prompt-based evaluation techniques show promise [13, 14], but their sensitivity to factual perturbations in elaborative content remains limited [12].
The lack of suitable benchmark datasets further hinders progress. Many existing datasets are constructed from LLM-generated summaries or apply rule-based perturbations to simulate non-factual content. For example, FactPICO provides expert annotations for plain language summaries of randomized controlled trial abstracts, focusing on PICO elements and evidence inference [9]. However, it includes factuality labels only for added content, leaving simplified sentences unannotated, which are generated by LLMs and potentially inaccurate. In contrast, APPLS perturbs human-written summaries using rule-based transformations [12], but cannot ensure that the resulting outputs remain coherent or factually plausible. These limitations underscore the need for high-quality, sentence-level annotations grounded in human-authored summaries.
Figure 1: Overview of PlainQAFact. A fine-tuned classifier first identifies the sentence type, involving either source simplification or elaborative explanation. Then, a QA-based evaluation pipeline performs answer extraction, question generation, question answering, and answer overlap evaluation. For elaborative content not present in the scientific abstract, PlainQAFact retrieves external knowledge to verify factual consistency. The illustrated example shows an elaborative explanation involving a “bronchodilator” not mentioned in the source abstract but verifiable through external evidence. PlainQAFact assigns a high score, reflecting strong alignment between the extracted and gold answers.
To address these challenges, we propose a targeted retrieval-based metric for factual consistency evaluation in PLS (Figure 1). We introduce a new expert-annotated dataset, PlainFact, consisting of human-written plain language summaries aligned with scientific abstracts. Each sentence is labeled with its type (elaborative explanation vs. source simplification), functional role, and alignment to the source (§3.1). Building on this, we present PlainQAFact, a dual-stage QA-based evaluation metric that selectively applies retrieval only to elaborative explanations (§3.2). This design ensures both efficiency and fidelity in evaluating factual consistency. Experiment results on several PLS datasets demonstrate the effectiveness of our approach, particularly in assessing complex, elaborative content (4, §5). Taken together, our work presents the first evaluation metric designed for PLS factual consistency evaluation, providing the community with both a robust benchmark and a practical tool to advance reliable and safe plain language communication in the medical domain.
@@ 2 Related Work
The primary approach for evaluating plain language generation combines automated metrics with human evaluation [3, 10]. While human assessment provides a thorough analysis [11], its high cost and time demands make it impractical for large-scale datasets. Evaluating factual consistency in biomedical PLS is particularly challenging, as it requires domain expertise. Entailment- [15, 16], similarity- [17, 18], model- [19] and QA-based [20, 21] metrics are commonly used for factual consistency assessment but rely heavily on high-quality reference summaries, which are often unavailable or difficult to obtain for PLS. Recent advancements in prompt-based evaluation show promise [13]; however, their sensitivity to factuality perturbations in PLS remains limited [12]. To address these limitations, we propose a reference-free solution for factual consistency evaluation of PLS that effectively assesses factual consistency with external information retrieval to augment the reference summary.
@@
Retrieval-augmented methods enhance text generation by extracting relevant information from external sources to supplement input queries [22]. These methods have been shown to be effective in open-domain QA [23, 24], knowledge-based QA [25], and multi-step reasoning [26, 27]. In the context of PLS, retrieval from structured knowledge bases (KBs) has been shown to improve factual accuracy compared to language models alone [28]. However, retrieval-augmented approaches have not been extensively explored for factual consistency evaluation in PLS, despite their potential for addressing elaborative explanations. In this work, we investigate retrieval-augmented QA to enhance PLS factual consistency assessment while also examining its limitations.
3 Methods
We first introduce the protocol of PlainFact dataset curation (§3.1). Based on this dataset, we propose PlainQAFact, a two-stage retrieval-augmented QA framework for factual consistency evaluation in PLS tasks (§3.2).
3.1 PlainFact Benchmark
To develop a high-quality factual consistency evaluation benchmark in PLS tasks, we collect a subset from the largest human-authored CELLS [28] dataset (§3.1.1) and hire domain experts to provide fine-grained sentence-level annotations (§3.1.2).
3.1.1 Human-Authored PLS Dataset
Rather than relying on LLM-generated plain language summaries, we construct our benchmark using human-authored summaries. CELLS [28], the largest parallel corpus of scientific abstracts and their corresponding plain language summaries, is written by the original authors and sourced from 12 biomedical journals. We primarily select data from the Cochrane Database of Systematic Reviews (CDSR) [^1] http:///www.cochranelibrary.com within CELLS, as CDSR contains systematic reviews that support evidence-based medical decision-making across healthcare domains [29]. Since systematic reviews represent the highest level of scientific evidence, this selection enhances the factual rigor of our dataset. To ensure readability, we filter the 200 most readable plain language summaries based on average scores from three standard readability metrics: Flesch-Kincaid Grade Level (FKGL) [30], Dale-Chall Readability Score (DCRS) [31], and Coleman-Liau Index (CLI) [32]. Additionally, given the collected plain language summaries from CELLS are all factual, we further conduct sentence-level perturbation to transform each plain language sentence into incorrect ones. Specifically, we use GPT-4o [^2] We use gpt-4o-2024-11-20 version for all experiments of GPT-4o in this paper. to perturb each plain language sentence based on prompts introduced in Guo et al. [12]. Therefore, for PlainFact, we have 400 summary-abstract pairs in total, where 200 factual pairs and 200 non-factual pairs.
Elaborative Explanation Source Simplification
@@ 3.1.1 Human-Authored PLS Dataset
FKGL $\downarrow$ 12.5 12.4
DCRS $\downarrow$ 11.3 11.6
CLI $\downarrow$ 13.5 13.9
@@
Table 1: Overview of the PlainFact benchmark. Medical experts annotated 200 pairs of plain language summaries and their corresponding scientific abstracts from the CELLS dataset [28], categorizing each plain language sentence as either a source simplification or an elaborative explanation related to the abstract. Lower scores of FKGL, DCRS, and CLI indicate better readability.
3.1.2 Expert Annotation
@@
Since each summary-abstract pair is authored by the same individual, we assume all information to be factual. The annotation aims to capture how plain language sentences relate to their scientific abstracts. Annotators analyze each plain language sentence across three dimensions: (1) Factuality type: identify whether a sentence is a source simplification (derived from the abstract) or an elaborative explanation (introducing new content); (2) Functional role: categorize the sentence as background, definition, example, method/result, or other; and (3) Sentence alignment: map each plain language sentence to its corresponding sentence(s) in the scientific abstract. Details of annotation guidelines are provided in A.
@@
Annotations are conducted by four independent annotators, each with at least a bachelor’s degree in biomedical sciences and prior fact-checking experience. Annotators are recruited via Upwork and compensated from \$15 to \$20 per hour. Each summary-abstract pair is annotated by two independent annotators, with disagreements resolved by a third. Inter-rater agreement, measured by Cohen’s Kappa for factuality type, functional role, and sentence alignment are 0.43, 0.60, and 0.55 respectively, indicated moderate agreement for all tasks [33].
@@
Table 1 summarizes the dataset characteristics. Notably, 44% of plain language sentences are elaborative explanations, highlighting their role in enhancing the readability of plain language summaries. 66% of these cannot be directly verified against the source abstract, which underscores the need for factuality evaluation methods that account for such phenomena. Moreover, elaborative explanations include more background, definitions, and examples than source simplifications. For annotation examples, see B.
3.2 PlainQAFact Framework
@@
PlainQAFact conducts fine-grained factual consistency evaluation for plain language summaries by first segmenting each summary into sentences, classifying as either simplification or explanation, and retrieving tailored external knowledge within a retrieval-augmented QA framework. Figure 1 provides an overview of its three key components: sentence-level classification (§3.2.1), domain knowledge retrieval (§3.2.2), and dual-stage QA-based factual consistency evaluation (§3.2.3).
3.2.1 Learned Factuality Type Classifier
@@
As elaborative explanations are prevalent in plain language generation and existing metrics struggle to capture added information [12], we first fine-tune a pre-trained language model to categorize factuality types of plain language sentences. Based on the factuality type annotations (source simplification vs. elaborative explanation) in PlainFact, we fine-tune the PubMedBERT-base model [^3] https://huggingface.co/microsoft/BiomedNLP-BiomedBERT-base-uncased-abstract-fulltext as a classifier. Our PlainFact is split 8:1:1 for training, validation, and testing. Additionally, we compare the pre-trained classifier with GPT-4o [34] as a zero-shot classifier (prompts are provided in C) to explore extending our factual consistency evaluation metric to domains and tasks lacking human-annotated data.
3.2.2 Domain Knowledge Retrieval
@@
Retrieval-augmented methods have proven effective for explanation generation [28, 35], making them a natural fit for evaluating elaborative explanations. Since our dataset is in the medical domain, we employ MedCPT [36], a retriever pre-trained on large-scale PubMed search logs to generate biomedical text embeddings. For external resources, we incorporate StatPearls [37] for clinical decision support and medical textbooks [38] for domain-specific knowledge.
3.2.3 QA Evaluation Components
@@
QA-based metrics have proved to be effective than other factual consistency evaluation metrics [12] in general summarization tasks and align more closely with human annotations [20, 39]. In this study, we adopt a QA-based approach as a backbone of our evaluation metric to verify factual consistency, incorporating answer extraction, question generation, question filtering, question answering, and answer overlap calculation modules.
@@
The first step in QA-based factual consistency evaluation is to extract answer entities (keyphrases) from plain language summaries as gold answer, then verify factuality by comparing them with answers generated by a QA model for the same questions. If the generated answer is correct or relevant, the summary is considered factual. The previous study used PromptRank [40], a keyphrase generation method based on the T5 model. To compare answer extraction strategies, we use PromptRank as the baseline, and evaluate LLM-based extractors, including an open-source Llama 3.1 8B Instruct model (Llama 3.1) [41] and close-source GPT-4o (GPT-4o) [34]. Prompts for Llama 3.1 and GPT-4o are in C.
@@
Given an input plain language summary, the QG model generates questions based on the extracted answers (§3.2.3) and the input plain language summary. Following previous studies [20, 21] [^4] These studies use allennlp, an open-source NLP research library built on PyTorch, which is no longer actively maintained., we fine-tune BART-large model on standard QG datasets, including SQuAD [42] and QA2D [43], for use as the QG model in our evaluation metric. The QG model will generate multiple relevant questions based on the input, which will increase the probability of verifying the factual consistency by answering these questions in the following stages.
@@
Questions generated by the QG model (§3.2.3) may not always be answerable. For example, the QG model generates “<What can occur over several days?>,” which is vague and hard to predict the correct answer. To prevent these unanswerable questions from impacting the final evaluation performance, we follow QAFactEval [39] and remove the unanswerable questions using a pre-trained Electra-large model [44]. During QF, the filtering model receives only the plain language summary and its corresponding questions. In the subsequent QA stage (§3.2.3), answers are extracted from the source for answer overlap evaluation, while QF only determines whether the questions can be answered by the plain language summary.
@@
The QA model extracts answers to answer the filtered questions from the source document. To prevent hallucinated output, we use an extractive QA model, a pre-trained Electra-large, which was the best performing QA model in QAFactEval [39].
@@
We evaluate the alignment between the gold and generated answers from the QA model using the baseline Learned Evaluation metric for Reading Comprehension (LERC) score [45] and BERTScore [46]. The final step in PlainQAFact is weighting the factual consistency scores for source simplification and elaborative explanation sentences. Specifically, the simplification score $s$ is computed using only abstracts as the source in QA, while the explanation score $e$ is calculated by incorporating both abstracts and retrieved knowledge as source for QA-based evaluation.
@@
where $n_{s}$ is the number of simplification sentences, and $n_{e}$ is the number labeled as explanation. $s_{\text{Avg.}}$ and $e_{\text{Avg.}}$ denote the average PlainQAFact scores computed for instances classified as simplification and explanation sentences, respectively.